# The Junction of Immersive Analytics and Virtual Reconstructions – A Case Study on the Mausoleum of Emperor Maxentius

### v1

@@ 5. Immersive Analytics for the Maxentius Mausoleum
This model visualises, as suggested by Johnson, a version of the Maxentius Mausoleum modelled closely on Tor de’ Schiavi (Johnson, 2009, p. 92). Consequently, model B can be considered a merger of models A and C. Thus, we have three reconstruction models of the same object (see Fig. 1), which mainly differ in the dome’s design and allow us to discuss which of the hypotheses about the part of the mausoleum is most plausible. We created the reconstructions with the 3D modelling software Blender. The components of the models consist of combined simple geometric forms, which were manipulated through manual node shifting to obtain necessary details. For each sub-component, once constructed from primitives, the measurements for those components were linearly scaled and applied. Thanks to the precise measurements we get from Rasch’s work, the models provide correct proportions, which helps us to obtain a fitting impression of the mausoleum’s former glory. In Fig. 3, we demonstrate this via overlaying Maxentius A transparently over a photogrammetry-based model of the mausoleum acquired from airborne imaging using a drone.
Figure 4. The software we provide together with our models. The panel, the miniature, and the actual visualisation can be seen.
[…]
An image from a user’s perspective, where the user sees an in-scale reconstruction of the Maxentius Mausoleum. What the user also sees is a miniature of the mausoleum reconstruction on a cube as well as a small panel with buttons with the names Maxentius A, B, and C, referring to the three available reconstructions.
@@
We provide software (see Fig. 4) that presents the three hypothesised models of the Maxentius mausoleum. The software provides a Virtual Reality environment, implemented with Unity C# and built to be used with an Oculus Quest 3 or Valve Index. Inside the immersive environment, the user will start in front of the stairway of Maxentius A. The user can move around, using the continuous movement metaphor and can snap-turn or move around freely, depending on the physical space available around them. When starting the software, in front of the user is a small panel with three buttons to switch between the three reconstruction models and a toggle to turn on or off a colouring akin to the extended matrix by Demetrescu et al. The buttons can be touched by the user with their hands. When the user touches one of the buttons, the presented model will switch, and the user will get haptic feedback.
To the right of the panel is a miniature of the model. Users can hold their virtual hands inside it and grab it with a simple grabbing motion. Upon grabbing, the user can control with their wrist the position of the sun and thus the light’s angle of incidence on the model. There is a small yellow sphere indicating the relative position of the light source, which imitates the sun so that users do not have to search for the real sun’s position. Once the user reaches for the miniature, a vibration of the controller indicates that they can now grab to rotate the light source. Once the grabbing motion is performed, a slightly more intense and more prolonged vibration pattern means that the wrist’s rotation now manipulates the light source’s angle. Furthermore, users can place small squares to leave small randomly coloured markers at points of interest behind. Lastly, the software can also be used without a VR headset with very simple keyboard and mouse controls, such that it can be viewed without the need of additional hardware. We provide the software via a Zenodo upload at (Kerle-Malcharek, 2025).
6. Discussion
Our exploration of IA applications in Virtual Archaeology argues for a significant potential for enhancing data understanding and analysis in archaeological reconstructions. While there exist works which advocate for the usefulness of immersive environments in Virtual Archaeology (Kyrlitsias et al., 2020), many works focus on dissemination or reconstruction. Therefore, this work yields significance for the field by presenting general considerations of how research can benefit from immersive technologies.
For our considerations, we utilise the Maxentius Mausoleum in Rome as a case study in the form of three digital reconstructions. The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea, which we also present, and are based on archaeological evidence and scholarly interpretations. We did this since we wanted to use reconstructions based on scientific evidence and valid assumptions to provide a sound foundation for our exploration.
Under the use of the models, we have created immersive reconstructions in the form of an interactive VR environment. This iVR environment serves as proof of concept of the potential of IA in archaeological analysis. We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation, offering new ways to interact with and comprehend complex spatiotemporal data.
@@ 6. Discussion
Our iVR environment allows users to manipulate lighting conditions, toggle between showing a colour coding for the certainty of single parts of the reconstruction and plain grey, place spatial markers and move around with a continuous movement metaphor. They use their hands to change the lighting and gain haptic feedback when using buttons inside the virtual space. We chose those interactions to align with IA research suggesting that embodied tools, among other concepts, can improve data comprehension (Marriott et al., 2018). We chose VR because the ability to interact with the reconstruction in an immersive environment supports a deeper understanding of spatial relationships. This is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy (Ragan et al., 2012). The colour coding we use builds upon established practices in Virtual Archaeology to represent reliability in reconstructions (Demetrescu, 2018).
Using methodologies from both fields shows that their respective strengths can be combined. The feature space of the respective topics in IA and Virtual Archaeology is vast, but we did not integrate a large number of potentially valuable features into our software. Our objective was not to provide comprehensive coverage but the demonstration of a number of possibilities as an example.
@@
The interaction possibilities we offer allow observers to gain a general impression of the reconstruction alternatives by switching through them on demand. Furthermore, observers can gain an intuition of the relative proportions of known parts of the mausoleum versus those which not known but can be derived from similar architecture. Changing the angle of incidence of the light source in combination with the miniature to see the relative position helps to gain an idea about lighting, especially inside the different interpretations of the mausoleum’s rotunda.
@@
The general implications of this work are significant for both archaeological methodology and public engagement. It emphasises a number of ways to interact with and analyse archaeological data in terms of representation and metadata and argues where future virtual archaeology projects could draw inspiration from. The ability to explore spatial relationships, lighting conditions, and levels of certainty in an immersive environment with embedded analysis information could lead to new research questions. Also, the presentation of data to people with and without domain background with the aspiration to make underlying reasoning understandable will benefit, as well. Scientifically, it emphasises a number of ways to interacting with and analysing archaeological data in terms of representation and metadata.
@@
Due to the character of this work, there are a number of limitations to consider. As for the models we created, as is usual in the domain, they also have parts that are more interpretative than others. The current implementation is focused on a single site and a specific set of interaction techniques. Although multi-modality is important, only haptic feedback and a visual representation of the models are available. Additionally, the use of VR hardware limits the accessibility for some users. While we do offer simple controls that allow to use it without VR hardware, the immersion factor that comes with VR is lost. Also, the considerations we make are a number of potential intersections of the two fields without taking deep considerations of how the design space of the different methods, like interaction design, would have to be applied.
@@
Thus, future work should include a variety of aspects. For one, a unified framework for creating immersive experiences would pose an asset. Such a framework should cover the inclusion and on-demand availability of historical background inside an immersive environment or the sonification of information, both of which have benefits. It may also be helpful to have ways of adjusting the light conditions tailored to the object’s geolocation and a chosen time and date. Another feature to further improve such software would be the simulation of crowds, as seen in the work on the Angkor Wat in Cambodia (Marriott et al., 2018, pp. 346–354).
@@
7. Conclusion
This work highlights Immersive Analytics’s potential for Virtual Archaeology, particularly for virtual reconstruction accessibility. We briefly discuss the possibilities of engaging representations of reconstructions and interactive representations of hypotheses. We exemplify the concept of integrating IA with our current research of the Maxentius Mausoleum, where these concepts will facilitate the analysis of the hypotheses regarding the construction and use of the Maxentius Mausoleum. To tackle the existing debates and derive new research questions, we provide an environment that showcases a few possibilities apart from standard visualisation. We virtualise the mausoleum’s reconstruction models and allow for some manipulation of the environment as well as changing and colourising the models akin to the Extended Matrix by Demetrescu et al. For example, we provide a general impression of how a light source must be positioned to obtain enough light inside the mausoleum, which may help rule out specific window configurations. Rasch’s proposed opening at the top seems to provide enough light to illuminate the whole room, while Johnson’s windows, as they are found at Tor de’ Schiavi, make it challenging to achieve that without additional light sources inside the mausoleum.
In our future research, we want to look into how the workflows from IA can be integrated into digital reconstruction. We want to combine existing modelling workflows for cultural heritage objects and advancements in data representation in immersive environments. We will continue to utilise the Maxentius Mausoleum for that and branch out to other objects, where appropriate. Overall, we present an exciting research direction that offers much potential for enhancing users’ experience of virtual archaeology environments.
8. Acknowledgements
We acknowledge funding by DFG, project ID 251654672 – TRR 161. We also acknowledge the support of Antonella Bidlingmaier, who contributed to the creation of the first iterations of the reconstructions of the Maxentius Mausoleum.

### v2

@@ IV Immersive Analytics for the Maxentius Mausoleum
Regarding the Maxentius Mausoleum, we chose Rasch’s work as a foundation for a first reconstruction (Maxentius A), which closely follows his suggested measurements. The second model (Maxentius B) assumes a closed dome following Johnson’s proposal [22, pp. 91-92] and adds seven round windows in the clerestory with a diameter of 1.80 meters considering the proportions of Tor de’ Schiavi [34, p. 28]. The third model (Maxentius C) adapts the flatter dome of Tor de’ Schiavi [34, p. 68] and the windows with a larger diameter. This model visualizes, as suggested by Johnson, a version of the Maxentius Mausoleum modeled closely on Tor de’ Schiavi [22, p. 92]. Consequently, model B can be considered a merger of models A and C. Thus, we have three reconstruction models of the same object (see Fig. 1), which mainly differ in the dome’s design and allow us to discuss which of the hypotheses about the part of the mausoleum is most plausible. We created the reconstructions with the 3D modeling software Blender. The components of the models consist of combined simple geometric forms, which were manipulated through manual node shifting to obtain necessary details. For each sub-component, once constructed from primitives, the measurements for those components were linearly scaled and applied. Thanks to the precise measurements we get from Rasch’s work, the models provide correct proportions, which helps us to obtain a fitting impression of the mausoleum’s former glory. In Fig. 3, we demonstrate this via overlaying Maxentius A transparently over a photogrammetry-based model of the mausoleum acquired from airborne imaging using a drone.
Figure 3: Cross-faded version of Maxentius A. Therefore, the different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
@@
We provide software (see Fig. 4) that presents the three hypothesized models of the Maxentius mausoleum. The software provides a VR environment, implemented with Unity C# and built to be used with an Oculus Quest 3 or Valve Index. We chose VR as it represents an s3D environment with head-tracking, which allows for improved spatial judgment. Inside the immersive environment, the user will start in front of the Maxentius A stairway. The user can move around, using the continuous movement metaphor, and can snap-turn or move around freely, depending on the physical space available around them.
Figure 4: The software we provide together with one of our models. The panel, the miniature, and the actual visualization can be seen.
There is a small panel with three buttons to switch between the three reconstruction models and a toggle to turn on or off a coloring akin to the extended matrix by Demetrescu by touching those buttons with their hands. Switching between models is especially relevant in a research context for comparing several different hypotheses directly while still in the iVR environment. Consequently, in dialogue with domain experts, compiling ways to compare the reconstruction models intuitively or conveniently helps to improve cognitive affordance.
To the right of the panel is a miniature of the model. By holding their virtual hands inside the miniature and performing a grabbing motion, a vibration of the controller indicates interaction, and they can control the light’s angle of incidence on the model. There is a small yellow sphere indicating the relative position of the light source at the miniature. This small sphere is intended to serve as a visualization to ease the intuition of where the light comes from.
This idiom is supposed to help when analyzing how and where sunlight comes from for the mausoleum. For our mausoleum scenario, it would make sense to look into relevant ways to interact with light sources. An optimized light interaction system could lead to a more fluent experience, which in turn may help experts justify the different shapes of the dome and the light conditions inside. Furthermore, users can place small squares to leave small randomly colored markers at points of interest behind. In general, different types of markers could be developed. For example, users could be enabled to colorize specific areas of the wall to mark the locations of the numerous drawings on the walls, which were added long after the mausoleum was originally built. The type of marker, color coding, or maybe even an acoustic signal could then again be investigated for its effectiveness in finding the different spots.
Lastly, the software can also be used without a VR headset with very simple keyboard and mouse controls, such that it can be viewed without the need for additional hardware. We provide the software via a Zenodo upload at [35].
V Discussion
With this work, we present three digital reconstructions of the Maxentius Mausoleum in Rome and utilize them to assess the value of IA for Virtual Archaeology. The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea and are based on archaeological evidence and scholarly interpretations. Such scientific evidence provides us with a sound foundation for our discussion in which we want to advocate for the use of immersive technologies more prominently for research purposes. While there exist works that point out the usefulness of immersive environments in Virtual Archaeology [29], they tend to focus on dissemination or reconstruction and not on how to improve the perception of a specific aspect, like the comparison of hypotheses.
@@ V Discussion
Using the three reconstruction models, we created immersive reconstructions in the form of an interactive VR (iVR) environment. This iVR environment serves as proof of concept of the potential of IA in archaeological analysis. We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation, offering new ways to interact with and comprehend complex spatio-temporal data. Our iVR environment allows users to manipulate lighting conditions, toggle between showing a color coding for the certainty of single parts of the reconstruction and plain grey, place spatial markers and move around with a continuous movement metaphor. They use their hands to change the lighting and gain haptic feedback when using buttons inside the virtual space. We chose those interactions to align with IA research suggesting that embodied tools, among other concepts, can improve data comprehension [11]. We chose VR because the ability to interact with the reconstruction in an immersive environment supports a deeper understanding of spatial relationships. This is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy [28]. The color coding we use builds upon established practices in Virtual Archaeology to represent reliability in reconstructions [9].
@@
As the feature space of IA topics and Virtual Archaeology is vast, we decided against integrating an exhaustive number of potential features into our software. Our objective was not to provide comprehensive coverage but to demonstrate a number of possibilities as an example. The interaction possibilities we offer allow observers to gain a general impression of the reconstruction alternatives by switching through them on demand. Furthermore, observers can gain an intuition of the relative proportions of known parts of the mausoleum versus those that are not known but can be derived from similar architecture. Through the miniature, users can see how the light source’s position changes relative to the reconstruction.
@@
The general implications of this work are significant for both archaeological methodology and public engagement. It emphasizes a number of ways to interact with and analyze archaeological data in terms of representation and metadata and argues where future virtual archaeology projects could draw inspiration from. The ability to explore spatial relationships, lighting conditions, and levels of certainty in an immersive environment with embedded analysis information could lead to new research questions.
@@
Due to the character of this work, there are a number of limitations to consider. As for the models we created, as is usual in the domain, they also have parts that are more interpretative than others. The current implementation is focused on a single site and a specific set of interaction techniques. Although multi-modality is important, only haptic feedback and a visual representation of the models are available. Additionally, the use of VR hardware limits the accessibility for some users. While we do offer simple controls that allow the software without VR hardware, the immersion factor that comes with VR is lost. Also, the considerations we make are a number of potential intersections of the two fields without taking deep considerations of how the design space of the different methods, like interaction design, would have to be applied.
@@
Thus, future work should include a variety of aspects. For one, a unified framework for creating immersive experiences would pose an asset. Such a framework should cover the inclusion and on-demand availability of historical background inside an immersive environment or the sonification of information, both of which have benefits. It may also be helpful to have ways of adjusting the light conditions tailored to the object’s geolocation and a chosen time and date.
@@
VI Conclusion
In this work, we present three reconstruction hypotheses of the Maxentius Mausoleum and the literature and typology that inspired them. In addition, we have created an immersive environment in which those reconstructions can be viewed and explored. The environment comes with a small number of integrated interaction and navigation possibilities to provide not only the models but also the means to explore them. Our second contribution is the exploration of the potential for Virtual Archaeology that Immersive Analytics as a research direction yields. We advocate the shift from ever-better representations, recording, and storage of cultural heritage items to tailored immersive experiences that improve accessibility and insight into existing data. The direction is to look at how virtual archaeology studies can be optimized by identifying the specific requirements of a given problem, like comparing different monument reconstruction interpretations and finding out how to support that. For argumentation, we outline the possibilities of engaging representations of reconstructions and interactive representations of hypotheses. We exemplify the concept of integrating IA with our current research on the Maxentius Mausoleum, where these concepts will facilitate the analysis of the hypotheses regarding its construction and use.
In our future research, we want to look into how the workflows from IA can be integrated into digital reconstruction. We like to combine existing modeling workflows for cultural heritage objects and advancements in data representation in immersive environments. We will continue to utilize the Maxentius Mausoleum for that and branch out to other objects, where appropriate. Overall, we present an exciting research direction that offers much potential for enhancing users’ experience of virtual archaeology environments.
Acknowledgment
We acknowledge funding by DFG, project ID 251654672–TRR 161 as well as the support of Antonella Bidlingmaier, who contributed to the creation of the first iterations of the reconstructions of the Maxentius Mausoleum.